OpenStudy (anonymous):

If you pour 100 grams of water from an 8 ounce can into a 16 ounce glass, by what amount does the density of the fluid change?

OpenStudy (aaronq):

density is defined as mass divided by the volume d=m/V so if you double the volume, the density will decrease by half d=1/2 double volume... d=1/4